# Formal Semantic Geometry over Transformer-based Variational AutoEncoder

### v2

@@ Abstract
Formal/symbolic semantics can provide canonical, rigid controllability and interpretability to sentence representations due to their localisation or composition property. How can we deliver such property to the current distributional sentence representations to control and interpret the generation of language models (LMs)? In this work, we theoretically frame the sentence semantics as the composition of semantic role - word content features and propose the formal semantic geometry. To inject such geometry into Transformer-based LMs (i.e. GPT2), we deploy Transformer-based Variational AutoEncoder with a supervision approach, where the sentence generation can be manipulated and explained over low-dimensional latent Gaussian space. In addition, we propose a new probing algorithm to guide the movement of sentence vectors over such geometry. Experimental results reveal that the formal semantic geometry can potentially deliver better control and interpretation to sentence generation.
@@ 1 Introduction
Formal semantics, which provides a canonical, granular, and rigid representation, have been investigated for thousands of years, such as Montague Semantics Dowty et al. (2012), Davidsonian Semantics Davidson (1967), Abstract Meaning Representation Banarescu et al. (2013), Semantic Role Labelling Palmer et al. (2010), and Argument Structure Theory (AST, Jackendoff (1992)). One typical characteristic of such formal semantics is the localisation or composition property. For example, in sentence: animals require oxygen for survival, the words are functionally combined into sentence semantics: $\lambda x(\text{animals}(x)\rightarrow\text{require}(x,\text{oxygen}))$ where $x$ is the variable of any entity within a logical structure. In this case, we can localise the sentence semantics by replacing $x$ with birds, etc. This localised process indicates the interpretation in Cognitive Science Smolensky (2006); Lees (1957). However, such localisation is precisely what current distributional semantics lack, thereby limiting their controllability and interpretability.
@@
Disentanglement Bengio (2013), which refers to the feature-dimension alignment (i.e., privileged basis Elhage et al. (2022)), can potentially provide such localisation, which has been widely investigated to localise image features, such as nose in facial images Esser et al. (2020); Jeon et al. (2019); Liu et al. (2021). In Transformers Vaswani et al. (2017), however, token embeddings, residual stream, and attention are non-privileged, meaning that multiple dimensions contribute to a feature. Although some prior studies explored the possibility of language disentanglement, most are focused on coarse-grained/task-specific semantic features, such as sentiment, within the context of style-transfer tasks John et al. (2019); Bao et al. (2019); Hu and Li (2021); Vasilakes et al. (2022); Gu et al. (2022); Liu et al. (2023a); Gu et al. (2023).
In this work, we focus on the localisation of general semantic features of sentences over distributional space to shorten the gap between deep latent semantics and formal linguistic representations Gildea and Jurafsky (2000); Banarescu et al. (2013); Mitchell (2023), integrating the flexibility of distributional-neural models with the properties of linguistically grounded representations, facilitating both interpretability and generative control from the perspective of formal semantics. We specifically choose the conceptual dense explanatory sentences from WorldTree Jansen et al. (2018) due to their clear formal semantic representation designed in the Explanatory Reasoning task.
@@
(1) We first propose a geometrical framework to present the formal semantic features of sentences as semantic role - word content pairs (denoted as role-content) from the perspective of AST Jackendoff (1992) within the compositional distributional model Clark et al. (2008). Subsequently, (2) we introduce a supervised approach for learning the role-content features of explanatory sentences in latent spaces. (3) Additionally, we propose a method to control sentence generation by navigating the sentence vectors across different role-content features within our geometric framework. (4) Our findings reveal that the role-content features are encoded as a convex cone in the latent sentence space (Figure 1). This semantic geometry facilitates the localisation of sentence generation by enabling the manipulation of sentence vectors through traversal and arithmetic operations within the latent space.
@@ 2 Related work
Integrating distributional semantics with formal / symbolic semantics is challenging in the field of artificial intelligence. In the Reasoning domain, for example, existing approaches usually perform symbolic behaviour via explicitly symbolic representation injection, including graph Khashabi et al. (2018); Khot et al. (2017); Jansen et al. (2017); Thayaparan et al. (2021), linear programming Valentino et al. (2022b); Thayaparan et al. (2024), adopting iterative methods, using sparse or dense encoding mechanisms Valentino et al. (2020); Lin et al. (2020); Valentino et al. (2022a); Bostrom et al. (2021), or synthetic natural language expression Clark et al. (2020); Yanaka et al. (2021); Fu and Frank (2024), among others. Comparatively, we explore the formal semantic property over distributional semantics via latent sentence geometry, which can potentially deliver better interpretation to current LMs.
@@
Disentanglement, refers to separating features along dimensions Bengio (2013), leading to clear geometric and linear representations. In the NLP domain, many studies explored the disentanglement between specific linguistic perspectives, such as sentiment-content John et al. (2019), semantic-syntax Bao et al. (2019), and negation-uncertainty Vasilakes et al. (2022), or syntactic-level disentanglement Mercatali and Freitas (2021); Felhi et al. (2022). However, a fundamental issue has been overlooked: the definition of disentanglement in the image domain Esser et al. (2020) cannot be directly applied to the context of computational linguistics due to the variability and complexity of language expression and high entanglement after current Transformer-based encoders. Therefore, we contribute to a new lens on the disentanglement (separation) of sentence features from the perspective of formal semantics.
@@ 3 Formal Semantic Geometry
In this section, we first define the sentence semantic features as semantic role - word content from the perspective of formal semantics. Then, we link the semantic features with distributional vector spaces. That is, each semantic role - word content is encoded as a convex cone in latent spaces.
@@
After defining the semantic features of sentences, we propose the concept of a convex cone of semantic feature. In linear algebra, a cone refers to a subset of a vector space that is convex if any $\alpha\overrightarrow{v_{i}}+\beta\overrightarrow{v_{j}}$ if any $\overrightarrow{v_{i}}$ and $\overrightarrow{v_{j}}$ belong to it. $\alpha$ and $\beta$ are positive scalars. Formally, the definition of convex cone, $C$, is described as a set of vectors: $C=\{x\in V|x=\sum_{i=1}^{n}\alpha_{i}v_{i},\alpha_{i}\geq 0,v_{i}\in R\}$ where $x$ is an element vector in vector space $\mathbb{R}$, $v_{i}$ are the basis vectors. $\alpha_{i}$ are non-negative scalars. In this context, we consider each role-content feature as a convex cone, $C$, corresponding to a hyperplane in high-dimensional vector space: $C_{c_{i},r_{i}}=\{t({c_{i}},{r_{i}})|t({c_{i}},{r_{i}})\in sem(s),s\in\textit{corpus}\}$ where $t({c_{i}},{r_{i}})$ represents the basis vector in $C_{c_{i},r_{i}}$ (Figure 2). According to set theory, we can define the formal semantic space as follows:
@@ 4 Geometrical Formal Semantic Control
Arithmetic has been considered a common way to control word or sentence semantics over latent spaces Mikolov et al. (2013b). E.g., the addition operation can steer the sentence semantics Shen et al. (2020); Mercatali and Freitas (2021); Liu et al. (2023b), or linear interpolation can generate smooth intermediate sentences Hu et al. (2022). However, they lack an explanation for these phenomena. In this section, we show that our geometrical framework can provide an intuitive explanation for these phenomena.
For linear interpolation, for example, it takes two sentences $x_{1}$ and $x_{2}$ and obtains latent vectors $z_{1}$ and $z_{2}$, respectively. It interpolates a path $z_{t}=z_{1}\cdot(1-t)+z_{2}\cdot t$ with $t$ increased from 00 to $1$ by a step size of $0.1$. Given two sentences with one role-content set overlap, $C_{c_{j},r_{j}}$. We can describe:
@@
$=(1-t)\times sem(s_{1})+t\times sem(s_{2})$
=$=\{\{z_{1}\cdot(1-t)+z_{2}\cdot t\},\dots\{\dots\}\}\cap C^{s_{1(2)}}_{c_{j},r_{j}}$
@@
“Linear representation hypothesis” refers to high-level concepts being represented linearly as directions in representation space, which has been widely evaluated to interpret Large LMs’ mechanism Marks and Tegmark (2023); Xie et al. (2021); Wang et al. (2024); Jiang et al. (2024b); Park et al. (2023, 2024). However, a main challenge for this hypothesis is that it’s not clear what constitutes a “high-level concept”.
Our geometrical framework can further support and extend this hypothesis by answering what and how they are “linearly” encoded? For example, given a set of $N$ atomic sentences: $s_{i}$: bird is a kind of living thing varying the content of arg1. Their semantics can be described below:
$sem(s)=\{C^{s_{i}}_{c_{i},arg1},\dots\}\cap\dots\cap C_{livingthing,arg2}$
@@
Since we describe different sentence semantic features, $\{c_{i},r_{i}\}$, as distinct convex cones, $C_{c_{i},r_{i}}$, within a $N$-dimensional vector space, $V\in\mathbb{R}^{N}$, we can linearly divide each basis dimension, $i\in{\{1,\dots,N\}}$, into different value regions, $[a,b]^{(i)}$, based on minimal information entropy. Consequently, there is a sequence of dimensional subspaces for each semantic feature. Thus, movement between different $C_{c_{i},r_{i}}$ regions can be achieved by moving out the dimensional regions within this sequence. This process can be implemented via a decision tree, In figure 3, for example, we can move the sentence from $C_{pred,causes}$ to $C_{pred,means}$ by modifying the values started from dim 21 $\leq-0.035$, …, ending at dim 10 $\leq-1.11$. By traversing the tree path, we can control the sentence generation by moving between convex cones, detailed in Algorithm 1.
@@ 5 SRL-Conditional VAE
In this section, we investigate the architecture of VAE to integrate the latent sentence space with LMs and propose a supervision approach to learn defined semantic features (i.e., role-content).
@@
$P_{\theta}(x,y,z)=\underbrace{P_{\theta}(x|z,r)}_{likelihood}\times\underbrace{P_{\theta}(z|r)}_{prior}\times P(r)$
$Z$$X$$p_{\theta}(x|z)$
@@
Figure 4: Computational graph of generation stage, where left: standard VAE, right: CVAE.
Specifically, we use encoder (i.e., Bert) to learn the approximate posterior based on both semantic roles and tokens, and additionally, we separately inject the semantic roles into encoder to learn the prior distribution. Both semantic roles and latent variables are injected into the decoder to auto-encode the tokens. The CVAE is trained to maximize the conditional log-likelihood of $x$ given $r$, which involves an intractable marginalization over the latent variable $z$. Moreover, to avoid the KL vanishing problem, which refers to the Kullback-Leibler (KL) divergence term in the ELBO becomes very small or approaches zero, we select the cyclical schedule to increase weights of KL $\beta$ from 0 to 1 Fu et al. (2019) and a KL thresholding scheme Li et al. (2019) that chooses the maximum between KL and threshold $\lambda$. The final objective function can be described as follows: $\mathcal{L}_{\text{CVAE}}=-\mathbb{E}_{q_{\phi}(z|r,x)}\Big{[}\log p_{\theta}(x|z,r)\Big{]}+\beta\sum_{i}\max\left[\lambda,\text{KL}q_{\phi}(z_{i}|x,r)||p(z_{i}|r)\right]$ where $q_{\phi}$ represents the approximated posterior (i.e., encoder). $i$ is the $i$-th latent dimension.
@@ 6 Empirical analysis
In the experiment, we quantitatively and qualitatively evaluate the latent space geometry via 1.traversal, 2.arithmetic, and 3.guided traversal. All experimental details are provided in Appendix A.
@@ 6.1 Latent Traversal
Next, we employ t-SNE Van der Maaten and Hinton (2008) to statistically examine role-content features cluster and separation over latent space (i.e., natural clustering property Bengio (2013)). In the corpus, however, due to the small number of data points within each role-content cluster, t-SNE cannot capture the differences between clusters well, resulting in the visualized latent space not displaying good role-content separability (top in figure 5). Therefore, we increase the number of data points in different role-content clusters by traversing each and keeping those resulting data points with the same role-content. Then, we visualise the role-content cluster at the bottom of figure 5. We can find that the features are clustered and separated over the latent space. If this was not the case, after traversing the resulting vectors from the same role-content cluster, the visualization should show the same entanglement as the original datapoints distribution.
@@ 6.3 Guided Latent Traversal
Finally, we examine the latent space geometry with our algorithm 1. The categories mentioned next are chosen based on their frequencies to ensure the balance during the training of the classifier.
@@
Finally, we use classification metrics, including accuracy (separability) and recall (density), as proxy metrics to assess latent space geometry. As shown in Table 7, both predicate and argument1 show higher separation.

### v3

# Quasi-symbolic Semantic Geometry over Transformer-based Variational AutoEncoder

@@ Abstract
Formal/symbolic semantics can provide canonical, rigid controllability and interpretability to sentence representations due to their localisation or composition property. How can we deliver such property to the current distributional sentence representations to better control and interpret the generation of language models (LMs)? In this work, we theoretically frame the sentence semantics as the composition of semantic role - word content features and propose the formal semantic geometrical framework. To inject such geometry into Transformer-based LMs (i.e. GPT2), we deploy a supervised Transformer-based Variational AutoEncoder, where the sentence generation can be manipulated and explained over low-dimensional latent Gaussian space. In addition, we propose a new probing algorithm to guide the movement of sentence vectors over such geometry. Experimental results reveal that the formal semantic geometry can potentially deliver better control and interpretation to sentence generation.
@@ 1 Introduction
Formal semantics, which provides a canonical, granular, and rigid representation, have been investigated for thousands of years with well established theoretical frameworks, such as Montague Semantics Dowty et al. (2012), Davidsonian Semantics Davidson (1967), Semantic Role Labelling (SRL, Palmer et al. (2010)), and Argument Structure Theory (AST, Jackendoff (1992)). One typical characteristic of such formal semantics is the localisation or composition property. For example, in the sentence: animals require oxygen for survival, the words are functionally combined into sentence semantics: $\lambda x(\text{animals}(x)\rightarrow\text{require}(x,\text{oxygen}))$ where $x$ is the variable of any entity within a logical structure. In this case, we can localise the sentence semantics by replacing $x$ with birds, etc. This localised process indicates the interpretation in Cognitive Science Lees (1957); Smolensky (2006). However, such localisation is precisely what current distributional semantics lack, thereby limiting their controllability and interpretability.
@@
Disentanglement Bengio (2013), which refers to the feature-dimension alignment, can potentially provide such localisation, which has been widely investigated to localise image features, such as nose in facial images Esser et al. (2020); Jeon et al. (2019); Liu et al. (2021). In Transformers Vaswani et al. (2017), however, token embeddings, residual stream, and attention have the polysemanticity phenomenon Elhage et al. (2022), meaning that multiple dimensions contribute to a feature. Although some prior studies explored the possibility of language disentanglement, most are focused on coarse-grained/task-specific semantic features, such as sentiment, within the context of style-transfer tasks John et al. (2019); Bao et al. (2019); Hu and Li (2021); Vasilakes et al. (2022); Gu et al. (2022); Liu et al. (2023a); Gu et al. (2023).
In this work, we focus on the localisation of general semantic features of sentences over distributional space to shorten the gap between deep latent semantics and formal linguistic representations Gildea and Jurafsky (2000); Banarescu et al. (2013); Mitchell (2023), integrating the flexibility of distributional-neural models with the properties of linguistically grounded representations, facilitating both interpretability and generative control from the perspective of formal semantics. We specifically choose the conceptual dense explanatory sentences from WorldTree Jansen et al. (2018) due to their clear formal semantic representation designed in the explanatory, cognitive reasoning task.
@@
(1) We first propose a geometrical framework to present the formal semantic features of sentences as semantic role - word content pairs (denoted as role-content) from the perspective of AST Jackendoff (1992) within the compositional distributional model Clark et al. (2008). Subsequently, (2) we introduce a supervised approach for learning the role-content features of explanatory sentences in latent spaces. (3) Additionally, we contribute to a method to control sentence generation by navigating the sentence vectors across different role-content features within our geometric framework. (4) Our findings reveal that the role-content features are encoded as a convex cone in the latent sentence space (Figure 1). This semantic geometry facilitates the localisation of sentence generation by enabling the manipulation of sentence vectors through traversal and arithmetic operations within the latent space.
@@ 2 Related work
Integrating distributional semantics with formal / symbolic semantics is challenging due to the difficulty of optimisation over discrete space van Krieken et al. (2023). In the Reasoning domain, existing approaches usually perform symbolic behaviour via explicitly symbolic representation injection, including graph Khashabi et al. (2018); Khot et al. (2017); Jansen et al. (2017); Thayaparan et al. (2021), linear programming Valentino et al. (2022b); Thayaparan et al. (2024), adopting iterative methods, using sparse or dense encoding mechanisms Valentino et al. (2020); Lin et al. (2020); Valentino et al. (2022a); Bostrom et al. (2021), or synthetic natural language expression Clark et al. (2020); Yanaka et al. (2021); Fu and Frank (2024), among others. Comparatively, we explore the formal semantic property over distributional semantics via latent sentence geometry, which can potentially deliver better interpretation and control to current LMs.
@@
Disentanglement, refers to separating features along dimensions Bengio (2013), leading to clear geometric and linear representations. In the NLP domain, prior studies explored the disentanglement between specific linguistic perspectives, such as sentiment-content John et al. (2019), semantic-syntax Bao et al. (2019), and negation-uncertainty Vasilakes et al. (2022), or syntactic-level disentanglement Mercatali and Freitas (2021); Felhi et al. (2022). However, those approaches focused on disentangling coarse-grained/task-specific semantic features. In this work, we contribute to a new lens on the disentanglement (separation) of “general” sentence features from the perspective of formal semantics.
@@ 3 Formal Semantic Geometry
In this section, we first define the sentence semantic features as semantic role - word content from the perspective of formal semantics. Then, we link the semantic features with distributional vector spaces in which each semantic role - word content is encoded as a convex cone, as shown in Figure 1.
@@
After defining the semantic features of sentences, we propose the concept of a convex cone of semantic feature. In linear algebra, a cone refers to a subset of a vector space that is convex if any $\alpha\overrightarrow{v_{i}}+\beta\overrightarrow{v_{j}}$ if any $\overrightarrow{v_{i}}$ and $\overrightarrow{v_{j}}$ belong to it. $\alpha$ and $\beta$ are positive scalars. Formally, the definition of convex cone, $C$, is described as a set of vectors: $C=\{x\in V|x=\sum_{i=1}^{n}\alpha_{i}v_{i},\alpha_{i}\geq 0,v_{i}\in R\}$ where $x$ is an element vector in vector space $\mathbb{R}$, $v_{i}$ are the basis vectors. $\alpha_{i}$ are non-negative scalars. In this context, we consider each role-content feature as a convex cone, $C$, corresponding to a hypersolid in high-dimensional vector space: $C_{c_{i},r_{i}}=\{t({c_{i}},{r_{i}})|t({c_{i}},{r_{i}})\in sem(s),s\in\textit{corpus}\}$ where $t({c_{i}},{r_{i}})$ represents the basis vector in $C_{c_{i},r_{i}}$ (Figure 2). According to set theory, we can define the formal semantic space as follows:
@@ 4 Geometrical Formal Semantic Control
Arithmetic has been considered a common way to control word or sentence semantics over latent spaces Mikolov et al. (2013b). E.g., the addition operation can steer the sentence semantics Shen et al. (2020); Mercatali and Freitas (2021); Liu et al. (2023b), or linear interpolation can generate smooth intermediate sentences Hu et al. (2022). However, they lack an explanation for these phenomena. We show that our geometrical framework can provide an intuitive explanation for these phenomena.
For linear interpolation, for example, it takes two sentences $x_{1}$ and $x_{2}$ and obtains latent vectors $z_{1}$ and $z_{2}$, respectively. It interpolates a path $z_{k}=z_{1}\cdot(1-k)+z_{2}\cdot k$ with $k$ increased from 00 to $1$ by a step size of $0.1$. Given two sentences with one role-content set overlap, $C_{c_{j},r_{j}}$. We can describe:
@@
$=(1-k)\times sem(s_{1})+k\times sem(s_{2})$
=$=\{\{z_{1}\cdot(1-k)+z_{2}\cdot k\},\dots\{\dots\}\}\cap C^{s_{1(2)}}_{c_{j},r_{j}}$
@@
“Linear representation hypothesis” refers to high-level concepts being represented linearly as directions in representation space, which has been widely evaluated to interpret Large LMs’ mechanism Marks and Tegmark (2023); Xie et al. (2021); Wang et al. (2024); Jiang et al. (2024b); Park et al. (2023, 2024). However, a main challenge for this hypothesis is that it’s not clear what constitutes a high-level concept.
Our geometrical framework can further support and extend this hypothesis by answering the questions: What and how are they “linearly” encoded? For example, given a set of $N$ atomic sentences: $s_{i}$: bird is a kind of living thing varying the content of arg1. Their semantics can be described below:
$sem(s)=\{C^{s_{i}}_{c_{i},arg1},\dots\}\cap\dots\cap C_{living~{}thing,arg2}$
@@
Since we describe different sentence semantic features, $\{c_{i},r_{i}\}$, as distinct convex cones, $C_{c_{i},r_{i}}$, within a $N$-dimensional vector space, $V\in\mathbb{R}^{N}$, we can linearly divide each basis dimension, $i\in{\{1,\dots,N\}}$, into different value regions, $[a,b]^{(i)}$, based on minimal information entropy. Consequently, there is a sequence of dimensional subspaces for each semantic feature. Thus, movement between different $C_{c_{i},r_{i}}$ regions can be achieved by moving out the dimensional regions within this sequence. This process can be implemented via a decision tree. In figure 3, for example, we can move the sentence from $C_{pred,causes}$ to $C_{pred,means}$ by modifying the values started from dim 21 $\leq-0.035$, …, ending at dim 10 $\leq-1.11$. By traversing the tree path, we can control the sentence generation by moving between convex cones, detailed in Algorithm 1.
@@ 5 SRL-Conditional VAE
In this section, we investigate the architecture of VAE to integrate the latent sentence space with LMs and propose a supervision approach to learn formal semantic geometry (i.e., role-content).
@@
$P_{\theta}(x,r,z)=\underbrace{P_{\theta}(x|z,r)}_{likelihood}\times\underbrace{P_{\theta}(z|r)}_{prior}\times P(r)$
pregrouptokenslanguagecomposetype logic
(a) CDM Clark et al. (2008)
@@
(b) SRL-Conditional VAE
Figure 4: Comparison between Compositional Distributional Model (CDM) (left) and SRL-Conditional VAE (right).
Specifically, we first model the categorical structures by encoding the semantic roles sequence to learn the prior distribution with parameters $\mu^{(srl)}$ and $\sigma^{(srl)}$. Then, we jointly encode semantic roles and lexical tokens to learn the approximate posterior parameterised by $\mu$ and $\sigma$. By minimising the Kullback-Leibler (KL) divergence between prior and approximate posterior, the semantic features can be encoded in the latent sentence space. Moreover, to avoid the KL vanishing problem, which refers to the KL term in the ELBO becomes very small or approaching zero, we select the cyclical schedule to increase weights of KL $\beta$ from 0 to 1 Fu et al. (2019) and a KL thresholding scheme Li et al. (2019) that chooses the maximum between KL and threshold $\lambda$. The final objective function can be described as follows:
$\mathcal{L}_{\text{CVAE}}=$ $-\mathbb{E}_{q_{\phi}(z|r,x)}\Big{[}\log p_{\theta}(x|z,r)\Big{]}$
$+\beta\sum_{i}\max\left[\lambda,\text{KL}q_{\phi}(z_{i}|x,r)||p(z_{i}|r)\right]$
where $q_{\phi}$ represents the approximated posterior (i.e., encoder). $i$ is the $i$-th latent dimension.
@@ 6 Empirical analysis
In the experiment, we quantitatively and qualitatively evaluate the latent space geometry via geometrical probing approaches: (1) traversal, (2) arithmetic, and (3) guided traversal. All experimental details are provided in Appendix A.
@@ 6.1 Latent Traversal
Next, we employ t-SNE Van der Maaten and Hinton (2008) to examine role-content features cluster and separation over latent space (i.e., natural clustering property Bengio (2013)). In the corpus, however, due to the small number of data points within each role-content cluster, t-SNE cannot capture the differences between clusters well, resulting in the visualized latent space not displaying good role-content separability (top in figure 5). Therefore, we increase the number of data points in different role-content clusters by traversing each and keeping those resulting data points with the same role-content. Then, we visualise the role-content cluster at the bottom of figure 5. We can find that the features are clustered and separated over the latent space. If this was not the case, after traversing the resulting vectors from the same role-content cluster, the visualization should show the same entanglement as the original datapoints distribution.
@@ 6.3 Guided Latent Traversal
Finally, we examine the semantic geometry via algorithm1. The categories selected below are chosen based on their frequencies, ensuring a balanced distribution during the classifier’s training process.
@@
Finally, we use classification metrics, including accuracy (separability) and recall (density), as proxy metrics to assess latent space geometry. As shown in Table 7, all features show higher separation where argument1 leads to the highest separation, indicating latent space geometry.